A dog and a rabbit are 180 meters apart The dog chases the rabbit. For every 9 meters that the dog runs, the rabbit jumps 5 meters. The distance, in meters, that the dog must run in order to overtake the rabbit is
(a) 184 (b) 189 (c) 225 (d) 405 (e) 1620

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

let them meet after x+180
The time taken by dog to travel 180+x = time taken by rabbit to go x meters
time = distance / speed
x+180 / 9= x/5
4x+900=9x
9x-5x=900
x=900 /4
225
225+180= 405 meters

Answer by Alan3354(69443)

A dog and a rabbit are 180 meters apart The dog chases the rabbit. For every 9 meters that the dog runs, the rabbit jumps 5 meters. The distance, in meters, that the dog must run in order to overtake the rabbit is
(a) 184 (b) 189 (c) 225 (d) 405 (e) 1620
----------------------
The dog gains on the rabbit at 4 meters/unit time (whatever the units are)
180/4 = 45 units
9*45 = 405 meters
d